OpenStudy (anonymous):

what is the height of a triangle with the area of 64

OpenStudy (anonymous):

Let h = height, then b = (1/2)h A = (1/2)bh 64 = (1/2)(1/2)h*h 64*4 = h^2 256 = h^2 h = 16

OpenStudy (anonymous):

is you a majic majic ninja......WHAT!

OpenStudy (anonymous):

\(\cal\Huge\color{blue}{Welcome~}\color{aqua}{To~}\color{purple}{OpenStudy!}\)

OpenStudy (cwrw238):

Note: this answer is correct only if the base is half the height - this was not indicated in the question